# PFT: Phonon Fine-tuning for Machine Learned Interatomic Potentials

### v2

@@ Abstract
Many materials properties depend on higher-order derivatives of the potential energy surface, yet machine learned interatomic potentials (MLIPs) trained with standard a standard loss on energy, force, and stress errors can exhibit error in curvature, degrading the prediction of vibrational properties. We introduce phonon fine-tuning (PFT), which directly supervises second-order force constants of materials by matching MLIP energy Hessians to DFT-computed force constants from finite displacement phonon calculations. To scale to large supercells, PFT stochastically samples Hessian columns and computes the loss with a single Hessian-vector product. We also use a simple co-training scheme to incorporate upstream data to mitigate catastrophic forgetting. On the MDR Phonon benchmark, PFT improves Nequix MP (trained on Materials Project) by 55% on average across phonon thermodynamic properties and achieves state-of-the-art performance among models trained on Materials Project trajectories. PFT also generalizes to improve properties beyond second-derivatives, improving thermal conductivity predictions that rely on third-order derivatives of the potential energy.
@@ 1 Introduction
We evaluate PFT by fine-tuning the Nequix MP foundation model (Koker et al., 2025) and benchmarking on held out calculations from the PBE MDR Phonon benchmark (Togo, 2023b; Loew et al., 2025). At a fraction of the cost of the initial pretraining, PFT reduces property error by 55% on average across maximum phonon frequency, vibrational entropy, Helmholtz free energy, and heat capacity, achieving state-of-the-art performance among models trained on MPtrj (Deng et al., 2023; Jain et al., 2013). Furthermore, we demonstrate that PFT improves generalization to third-order derivatives, improving Matbench Discovery thermal conductivity predictions from 0.446 to 0.306 $\kappa_{\text{SRME}}$, which is also state-of-the-art among MPtrj-trained models. Lastly, we demonstrate that through co-training, PFT degrades performance by less than 1% on the Matbench Discovery stability classification task, which is otherwise greatly impacted by training on phonon data alone.
@@
  • A simple co-training recipe that mitigates catastrophic forgetting by interleaving pre-training data into the finetuning procedure.
@@ 2.2 Harmonic phonons and vibrational properties
where $i$ and $j$ are Cartesian indices, and $a$ and $b$ are atom indices.
@@ 2.3 Phonon calculations
Phonon calculations are usually conducted with DFT in one of two ways: 1) using density functional perturbations theory (DFPT), or 2) DFT and with finite displacements. The latter is more common due to its generality across commonly used functionals (Miranda, ).
@@
$F_{b,j}(\Delta r_{a,i})$ are forces where atom $a$ is displaced in direction $i$, and generally $F_{b,j}=0$ due to the atom positions being at equilibrium. This approximation works well due to the smoothness of DFT under small displacements, commonly 0.01 Å for these calculations. The number of finite-displacement calculations (3N) can typically be significantly reduced by leveraging the symmetry of the crystal (Togo et al., 2023)
@@
MLIPs can be used for phonon calculations in the same way as DFT, constructing the force constants by conducting force calculations under finite displacement (Fig. 1c(i-ii)), which has shown success in prediction of both second-order (Loew et al., 2025) and third-order (Póta et al., 2024) phonon properties.
@@ 3.1 Hessian error and vibrational properties
Because phonon spectra and derived vibrational properties are functions of the second-order force constants, models must match the curvature of the PES to achieve accurate phonon predictions. Using the finite displacement method above, we calculate the error in the second-order force constants, or Hessians of several models trained on MPtrj (Deng et al., 2023; Jain et al., 2013), evaluating them on a portion of the force constants and phonon properties from the PBE MDR phonon dataset (Loew et al., 2025; Togo, 2023b). Figure 2 shows that lower hessian error correlates with lower error on multiple thermodynamic properties. This motivates that training directly on second-order force constants results in more accurate predictions of these properties.
@@ 3.2 Phonon fine-tuning (PFT)
We propose phonon fine-tuning (PFT), a method for finetuning MLIPs directly on DFT-computed force constants. $\mathcal{L}_{\text{PFT}}$ is a four term loss function, with terms that minimize error in the energy, force, stress, and second-order force constants:
@@
where and $N_{a}$ is the number of atoms in a system. Note that again the batch dimension is omitted for readability. $\hat{E}$ is the predicted potential energy from the neural network; the first three terms follow the standard EFS loss from Eq. (3) with an MAE loss on energy and stress and a $\ell_{2}$ loss on forces. Forces are predicted as the negative derivative of energy with respect to position $\mathbf{r}$, and stress is predicted as the derivative of energy with respect to a symmetric strain tensor. The second-order force constants are predicted analytically as the hessian of the energy with respect to two atom positions.
@@
Since the DFT phonon calculations themselves consist of energy, force, and stress calculations under displacements of atoms, it is reasonable to assume that one could simply fine-tune models directly on this displacement data, as it contains sufficient information for constructing the full Hessian. Empirically we find this not to be the case, and observe that directly finetuning on the phonon displacements results in a worse hessian error. This suggests that the standard practice of EFS training on DFT calculations of rattled, or perturbed structures (Barroso-Luque et al., 2024; Kaplan et al., 2025) may not be sufficient for correctly modeling the PES curvature.
Figure 3: Training ablation. The top row shows energy, force, and stress errors on the MPtrj validation set, while the bottom row shows energy, force, stress, and hessian errors on the MDR Phonon validation set during finetuning of Nequix MP. We compare phonon fine-tuning with and without co-training on MPtrj, as well directly finetuning energy, force, and stress on the phonon displacement calculations. Co-training mostly mitigates degradation of MPtrj performance at only a slight increase in hessian error, and that training directly on displacements worsens hessian performance over the base model. *We note that larger energy errors on MDR Phonon data are likely due to a mismatch in energies between MPtrj and MDR Phonon, see Sec. A.1.
@@ 3.3 Efficient computation via Hessian-vector product
which can be computed using a Hessian-vector product (HVP) without materializing the full hessian using a forward-mode Jacobian-vector product (JVP) through a reverse-mode gradient (Pearlmutter, 1994). In practice, Hessian columns across a full batch of structures is computed by forming a single graph of atom positions where each structure is not connected, concatenating the sampled $\mathbf{v}$ for each structure into a single vector, and computing the HVP with respect to the sum of energies across all structures. This facilitates calculation of the loss a with single HVP call, and enables training on GPUs that would otherwise not have enough memory for a single Hessian calculation, especially for the large supercells that are needed for accurate phonon calculations. Note that during training, this results in a “triple-backward” as gradient-based optimization requires a derivative of the HVP with respect to model weights.
@@ 3.4 Co-training
While fine-tuning models on phonon data will more closely align the curvature of the PES with that of DFT, this procedure may lead to catastrophic forgetting (French, 1999), where the fine-tuning procedure effects the performance of the model on its original upstream training data. This may be an issue as phonon calculations are always done at equilibrium, which can could cause forgetting for out-of-distribution non-equilibrium configurations. Furthermore, if the relative quantity and diversity of the phonon fine-tuning dataset is less than that of the original pretraining or fine-tuning datasets (such as the case in our experimental setting)
@@ 3.5 Analytical property prediction
As discussed, most existing methods for evaluating MLIPs on higher order derivatives of the PES do so with finite displacement (Loew et al., 2025; Póta et al., 2024). This is typically accurate if the smoothness and continuity of the PES is taken into account when designing the neural network architecture; however, it does introduce an additional hyper-parameter in the form of displacement distance, which can subtly effect results (Fu et al., 2025).
@@ 4.1 Training
Figure 4: Phonon band structures. For ease of visualization, we display the phonon band structure for the three materials in the test split of MDR Phonon with the fewest number of atoms. In the case of Carbon structure (mp-990448), it results in a dynamically stable structure similar to DFT while the non-PFT model shows imaginary mode around the $q$-point AA. Overall, we find that that PFT generally produces bands with closer alignment to those of the DFT ground truth. More phonon bandstructures are provided in Fig. A.2.
For phonon calculation data $\mathcal{D}_{\text{phonon}}$, we use MDR Phonon calculation database (Togo, 2023b), which was recalculated by (Loew et al., 2025) using the PBE exchange correlation functional in order to match the settings of the Materials Project data. These calculations were conducted using finite displacement (see Sec. 2.3), and contain the original energy/force/stress calculations at each displacement. While the phonon data itself does not contain force, or stress labels needed for the PFT loss function (Eq. 8), we assume the force and stress labels are zero due to the strict structural relaxation procedure done prior to the phonon calculation. The energy label is computed by multiplying the provided unitcell energy by the number of repetitions within the supercell used for calculations. Of the 9,959 materials within the dataset, all but 91 exist in MPtrj based on mp-id. These 91 materials along with randomly selected calculations from the remaining data are used as a test set of 1,000 materials, with the rest of the data split into training/validation subsets with proportion 95/5. This ensures any benefit we see from PFT is due to the additional phonon information, and not novel chemistry or geometries. We do note there is a discrepancy between the energies of materials in phonon data and MPtrj, which we detail in Sec. A.1.
We perform PFT both with ($K=4$) and without ($K=0$) co-training for 200 epochs on the MDR Phonon data. With co-training, this requires about 140 A100 hours, with about 60 A100 hours for PFT without co-training — significantly fewer than the 500 A100 hours needed to train the base Nequix MP model and a small fraction of other competitive MPtrj-trained MLIPs (Koker et al., 2025). For details on other hyper-parameters and their selection see Sec. A.2.
@@ 4.2 Phonon properties
Table 1: Evaluation of models on held-out MDR Phonon data. Metrics are MAE of maximum phonon frequency $\omega_{\text{max}}$ (K), vibrational entropy $S$ (J/K/mol), Helmholtz free energy $F$ (kJ/mol) and heat capacity at constant volume $C_{V}$ (J/K/mol).
@@
Table 1 shows the errors of each model on phonon related properties maximum phonon frequency, vibrational entropy, Helmholtz free energy, and heat capacity at constant volume. From the Nequix MP base model, PFT results in an average reduction in property prediction error of 55%, and acheives state-of-the-art error across all metrics for models trained on MPtrj materials despite being the smallest model. We find no significant differences between properties computed from forces constants obtained via finite displacement or AD, although the PFT model without co-training demonstrates a slight improvement in prediction error.
@@ 4.3 Thermal conductivity
To demonstrate the effect of PFT on properties that are higher-order derivatives of the PES, we evaluate the model on thermal conductivity, a task from the Matbench Discovery (Riebesell et al., 2025) benchmark. Thermal conductivity $\kappa$ is a function of third-order force constants (Póta et al., 2024), or the third derivative of the potential energy. Just as with second order force constants, third order force constants can be calculated with finite displacement, leveraging symmetry to reduce the number of computations (Togo, 2023a).
Table 2: Matbench Discovery “compliant” leaderboard for thermal conductivity, measured in symmetric relative mean error in predicted phonon mode contributions to thermal conductivity $\kappa_{\mathrm{SRME}}$.
@@
We follow the procedure used by Matbench Discovery (Riebesell et al., 2025; Póta et al., 2024) performing a structural relaxation and then calculation of the third-order force constants with a displacement of 0.03 Å. Results from all of the other MPtrj-trained, or “compliant” models (Neumann et al., 2024; Barroso-Luque et al., 2024; Zhou et al., 2025; Batatia et al., 2025a; Zhang et al., 2025; Yan et al., 2025; Park et al., 2024; Bochkarev et al., 2024; Tan et al., 2025; Koker et al., 2025; Chen et al., 2025; Fu et al., 2025) sourced from the Matbench Discovery leaderboard are shown in Table 2.
Similarly to phonon properties we find a reduction in error of 31% from the base Nequix MP model, and state-of-the-art performance for MPtrj-trained models. In Fig. 6, we show macroscopic conductivity predictions for two zinc blends, AgI and BeO, showing again that PFT aligns predictions closer to the ground truth DFT calculation. This shows that PFT not only improves second order vibrational properties, but also generalizes to third order, or anharmonic, vibrational properties.
@@ 4.4 Matbench Discovery
The main Matbench Discovery task (Riebesell et al., 2025) consists of a geometry optimization followed by energy prediction to determine material stability. Performance on this task is less influenced by curvature of the PES and largely dictated by energy error so it highlights any degradation of energy prediction caused by finetuning. Table 3 shows the evaluation of the geometry optimization task, measured in in root mean squared distance (RMSD) from the DFT-optimized structures as well as the F1 score of the stability classification task.
Table 3: Evaluation of models on Matbench Discovery, consisting of geometry optimization, measured in RMSD (Å) and stable/unstable material classification F1 on unique prototypes.
@@
We find that neither PFT models improve performance over the base model. This may be expected, especially for F1, as the finetuning data has no additional energy data from extra chemistries or geometries, so performance of the architecture in this task will be saturated by the original MPtrj training data. However, the co-training is vital for preserving performance on this task; while the non co-trained PFT model worsens RMSD and F1 by 5% and 60% respectively, co-training reduces this difference to less than a 2% increase in geometry optimization RMSD and less than 1% reduction in stability classification F1.

### v3

@@ Abstract
Many materials properties depend on higher-order derivatives of the potential energy surface, yet machine learned interatomic potentials (MLIPs) trained with a standard loss on energy, force, and stress errors can exhibit error in curvature, degrading the prediction of vibrational properties. We introduce phonon fine-tuning (PFT), which directly supervises second-order force constants of materials by matching MLIP energy Hessians to DFT-computed force constants from finite displacement phonon calculations. To scale to large supercells, PFT stochastically samples Hessian columns and computes the loss with a single Hessian-vector product. We also use a simple co-training scheme to incorporate upstream data to mitigate catastrophic forgetting. On the MDR Phonon benchmark, PFT improves Nequix MP by 55% on average across phonon thermodynamic properties and achieves state-of-the-art accuracy among models trained on Materials Project trajectories. PFT also generalizes to improve properties beyond second-derivatives, improving thermal conductivity predictions that rely on third-order derivatives of the potential energy.
@@ 1 Introduction
We evaluate PFT by fine-tuning the Nequix MP foundation model (Koker et al., 2025) and benchmarking on held out calculations from the PBE MDR Phonon benchmark (Togo, 2023b; Loew et al., 2025). At a fraction of the cost of the initial pretraining, PFT reduces property error by 55% on average across maximum phonon frequency, vibrational entropy, Helmholtz free energy, and heat capacity, achieving state-of-the-art accuracy among models trained on MPtrj (Deng et al., 2023; Jain et al., 2013). Furthermore, we demonstrate that PFT improves generalization to third-order derivatives, improving Matbench Discovery thermal conductivity predictions from 0.446 to 0.306 $\kappa_{\text{SRME}}$, which is also state-of-the-art among MPtrj-trained models. We demonstrate that through co-training, PFT degrades performance by less than 1% on the Matbench Discovery stability classification task, which is otherwise greatly impacted by training on phonon data alone. Lastly, we apply PFT to a stronger base model trained on OMat24 (Barroso-Luque et al., 2024), observing consistent benefits despite substantially more upstream data.
@@
  • A simple co-training recipe that mitigates catastrophic forgetting by interleaving pre-training data into the fine-tuning procedure.
@@ 2.2 Harmonic phonons and vibrational properties
where $i$, $j$ are Cartesian indices, and $a$, $b$ are atom indices.
@@ 2.3 Phonon calculations
Phonon calculations are usually conducted with DFT in one of two ways: 1) using density functional perturbation theory (DFPT), or 2) using DFT with finite displacements. The latter is more common due to its generality across commonly used functionals (Miranda, ).
@@
$F_{b,j}(\Delta r_{a,i})$ are forces where atom $a$ is displaced in direction $i$, and generally $F_{b,j}=0$ due to the atom positions being at equilibrium. This approximation works well due to the smoothness of DFT under small displacements, commonly 0.01 Å for these calculations. The number of finite-displacement calculations (3N) can typically be significantly reduced by leveraging the symmetry of the crystal (Togo et al., 2023).
@@
MLIPs can be used for phonon calculations in the same way as DFT, constructing force constants with force calculations under finite displacement (Fig. 1c(i-ii)), which has shown success in predicting second-order (Loew et al., 2025) and third-order (Póta et al., 2024) phonon properties.
@@ 3.1 Hessian error and vibrational properties
Because phonon spectra and derived vibrational properties are functions of the second-order force constants, models must match the curvature of the PES to achieve accurate phonon predictions. Using the finite displacement method above, we calculate the error in the second-order force constants, or Hessians of models trained on MPtrj (Deng et al., 2023; Jain et al., 2013), evaluating them on a portion of the force constants and phonon properties from the PBE MDR phonon dataset (Loew et al., 2025; Togo, 2023b). Figure 2 shows that lower Hessian error correlates with lower error on multiple thermodynamic properties. This motivates that training directly on second-order force constants results in more accurate predictions of these properties.
@@ 3.2 Phonon fine-tuning (PFT)
We propose phonon fine-tuning (PFT), a method for fine-tuning MLIPs directly on DFT-computed force constants. $\mathcal{L}_{\text{PFT}}$ is a four term loss function, with terms that minimize error in the energy, force, stress, and force constants:
@@
where $N_{a}$ is the number of atoms in a system. Note that again the batch dimension is omitted for readability. $\hat{E}$ is the predicted potential energy from the neural network; the first three terms follow the standard EFS loss from Eq. (3) with an MAE loss on energy and stress and a $\ell_{2}$ loss on forces. The second-order force constants are predicted analytically as the Hessian of the energy with respect to two atom positions.
@@
Since the DFT phonon calculations themselves consist of energy, force, and stress calculations under displacements of atoms, it is reasonable to assume that one could simply fine-tune models directly on this displacement data, as it contains sufficient information for constructing the full Hessian. Empirically we find this not to be the case, and observe that directly fine-tuning on the phonon displacements results in a significantly worse Hessian error (Fig. 3, Table 1). This suggests that the standard practice of EFS training on DFT calculations of rattled, or perturbed structures (Barroso-Luque et al., 2024; Kaplan et al., 2025) may not be sufficient for correctly modeling the PES curvature.
Figure 3: Training ablation. The left figure shows Hessian error on the MDR Phonon validation set, and the remaining figures show energy, force, and stress errors on the MPtrj validation set (top) and the MDR Phonon validation set (bottom). We compare phonon fine-tuning with and without co-training on MPtrj, as well as directly fine-tuning energy, force, and stress on the phonon displacement calculations. Co-training mostly mitigates degradation of MPtrj performance at only a slight increase in Hessian error. Training directly on displacements significantly worsens Hessian error over the base model. *We note that larger energy errors on MDR Phonon data are likely due to a mismatch in energies between MPtrj and MDR Phonon, see Sec. A.2.
@@ 3.3 Efficient computation via Hessian-vector product
which can be computed without materializing the full Hessian by using a Hessian-vector product (HVP), using a forward-mode Jacobian-vector product (JVP) through a reverse-mode gradient (Pearlmutter, 1994). In practice, Hessian columns across a full batch of structures are computed by forming a single graph of atom positions where each structure is not connected, concatenating the sampled $\mathbf{v}$ for each structure into a single vector, and computing the HVP with respect to the sum of energies across all structures. This facilitates calculation of the loss with a single HVP call, and enables training on GPUs that would otherwise not have enough memory for a single Hessian calculation, especially for the large supercells that are needed for accurate phonon calculations (see Sec. A.1). During training, this results in a “triple-backward” as gradient-based optimization requires a derivative of the HVP with respect to model weights.
@@ 3.4 Co-training
While fine-tuning models on phonon data will more closely align the curvature of the PES with that of DFT, this procedure may lead to catastrophic forgetting (French, 1999), where the fine-tuning procedure effects the performance of the model on its original upstream training data. This may be an issue as phonon calculations are always done at equilibrium, which can cause forgetting for out-of-distribution non-equilibrium configurations. Furthermore, the relative quantity and diversity of the phonon dataset may be less than the original pretraining or fine-tuning datasets (such as the case in our experimental setting).
@@ 3.5 Analytical property prediction
As discussed, most existing methods for evaluating MLIPs on higher order derivatives of the PES do so with finite displacement (Loew et al., 2025; Póta et al., 2024). This is typically accurate if the smoothness and continuity of the PES is taken into account when designing the neural network architecture; however, it does introduce an additional hyper-parameter in the form of displacement distance, which can subtly affect results (Fu et al., 2025).
@@
Figure 4: Phonon band structures. For ease of visualization, we display the phonon band structure for the three materials in the test split of MDR Phonon with the fewest number of atoms. In the case of Carbon structure (mp-990448), it results in a dynamically stable structure similar to DFT while the non-PFT model shows imaginary mode around the $q$-point AA. Overall, we find that PFT generally produces bands with closer alignment to those of the DFT ground truth. More phonon band structures are provided in Fig. A.3.
@@ 4.1 Training
For phonon calculation data $\mathcal{D}_{\text{phonon}}$, we use MDR Phonon calculation database (Togo, 2023b), which was recalculated by (Loew et al., 2025) using the PBE exchange correlation functional in order to match the settings of the Materials Project data. These calculations were conducted using finite displacement (see Sec. 2.3), and contain the original energy/force/stress calculations at each displacement. While the phonon data itself does not contain force, or stress labels needed for the PFT loss function (Eq. 8), we assume the force and stress labels are zero due to the strict structural relaxation procedure done prior to the phonon calculation. The energy label is computed by multiplying the provided unitcell energy by the number of repetitions within the supercell used for calculations. Of the 9,959 materials within the dataset, all but 91 exist in MPtrj based on mp-id. These 91 materials along with randomly selected calculations from the remaining data are used as a test set of 1,000 materials, with the rest of the data split into training/validation subsets with proportion 95/5. This ensures any benefit we see from PFT is due to the additional phonon information, and not novel chemistry or geometries. The training set then contains 8,510 structures derived from 301,414 displacement DFT calculations. We do note there is a discrepancy between the energies of materials in phonon data and MPtrj, which we detail in Sec. A.2.
We perform PFT both with ($K=4$) and without ($K=0$) co-training for 200 epochs on the MDR Phonon data. With co-training, this requires about 140 A100 hours, with about 60 A100 hours for PFT without co-training — significantly fewer than the 500 A100 hours needed to train the base Nequix MP model and a small fraction of other competitive MPtrj-trained MLIPs (Koker et al., 2025). To evaluate robustness to base model quality and scale of pre-training data, we pre-train a new base model, Nequix OAM (see Sec. A.3.1), and apply PFT using the same procedure as Nequix MP. For details on other hyper-parameters and their selection see Sec. A.3.
@@ 4.2 Phonon properties
Table 1: Evaluation of models on held-out MDR Phonon data. Metrics are MAE of maximum phonon frequency $\omega_{\text{max}}$ (K), vibrational entropy $S$ (J/K/mol), Helmholtz free energy $F$ (kJ/mol) and heat capacity at constant volume $C_{V}$ (J/K/mol). Additional upstream data (Sec. A.3.1).
@@
Table 1 shows the errors of each model on phonon related properties maximum phonon frequency, vibrational entropy, Helmholtz free energy, and heat capacity at constant volume. From the Nequix MP base model, PFT results in an average reduction in MAE of 55%, and achieves state-of-the-art error across all metrics for models trained on MPtrj materials despite being the smallest model. We find no significant differences between properties computed from forces constants obtained via finite displacement or AD, although the PFT model without co-training demonstrates a slight improvement in MAE. Applying PFT to Nequix OAM yields a similar average relative improvement of 50%. Notably, the MP PFT model exceeds the performance of the OAM base model, despite being trained with approximately two orders of magnitude fewer DFT calculations.
@@ 4.3 Thermal conductivity
To demonstrate the effect of PFT on properties that are higher-order derivatives of the PES, we evaluate the model on thermal conductivity, a task from the Matbench Discovery (Riebesell et al., 2025) benchmark. Thermal conductivity $\kappa$ is a function of third-order force constants (Póta et al., 2024), or the third derivative of the energy. Just as with second order force-constants, third order force constants can be calculated with finite displacement, leveraging symmetry to reduce the number of computations (Togo, 2023a).
Table 2: Matbench Discovery “compliant” leaderboard for thermal conductivity, measured in symmetric relative mean error in predicted phonon mode contributions to thermal conductivity $\kappa_{\mathrm{SRME}}$. Additional upstream data (Sec. A.3.1).
@@
We follow the procedure used by Matbench Discovery (Riebesell et al., 2025; Póta et al., 2024) performing a structural relaxation and then calculation of the third-order force constants with a displacement of 0.03 Å. Results from all of the other MPtrj-trained, or “compliant” models (Neumann et al., 2024; Barroso-Luque et al., 2024; Zhou et al., 2025; Batatia et al., 2025a; Zhang et al., 2025; Yan et al., 2025; Park et al., 2024; Bochkarev et al., 2024; Tan et al., 2025; Chen et al., 2025; Fu et al., 2025) sourced from the Matbench Discovery leaderboard are shown in Table 2.
Similarly to phonon properties we find a reduction in error of 31% from the base Nequix MP model, and state-of-the-art performance for MPtrj-trained models. For the OAM base model, PFT improves error by 26%. In Fig. 6, we show macroscopic conductivity predictions for two zinc blends, AgI and BeO, showing again that PFT aligns predictions closer to the ground truth DFT result. PFT not only improves second order vibrational properties, but also generalizes to third-order, or anharmonic, vibrational properties.
@@ 4.4 Matbench Discovery
The main Matbench Discovery task (Riebesell et al., 2025) consists of a geometry optimization followed by energy prediction to determine material stability. Performance on this task is less influenced by curvature of the PES and largely dictated by energy error, so it highlights any degradation of energy prediction caused by fine-tuning. Table 3 shows the evaluation of the geometry optimization task, measured in root mean squared distance (RMSD) from the DFT-optimized structures as well as the F1 score of the stability classification task.
Table 3: Evaluation of models on Matbench Discovery, consisting of geometry optimization, measured in RMSD (Å) and stable/unstable material classification F1 on unique prototypes. Additional upstream data (Sec. A.3.1).
@@
Nequix OAM 0.0659 0.874
Nequix OAM PFT 0.0666 0.864
We find that neither PFT models improve performance over the base model. This may be expected, especially for F1, as the fine-tuning data has no additional energy data from extra chemistries or geometries, so performance of the architecture in this task will be saturated by the original MPtrj training data. However, the co-training is vital for preserving performance on this task; while the non-co-trained PFT model worsens RMSD and F1 by 5% and 60% respectively, co-training reduces this difference to less than a 2% increase in geometry optimization RMSD and less than 1% reduction in stability classification F1.